OpenStudy (anonymous):

Which of the following is a solution of x2 + 2x = −2? x = 4 − 3i x = −4 + 3i x = −1 + i x = 1 − i

OpenStudy (kagıtucak):

x^2+2x+2=0 from the formula roots of the second order equations are (-b+- root(delta))/2a where b=2 and a=1 in this equation for delta the formula is b^2-4ac where c=2 so delta= 4-4*1*2=-4 for roots now (-2+- root(-4))/2 =(-2+-2i)/2 =-1+-i We have -1+i in options so thats the answer

OpenStudy (anonymous):

thank you!

OpenStudy (kagıtucak):

You're welcome!